OpenStudy (anonymous):

Can you show me step by step, I'm having trouble with these types of expressions. Simplify the expression. quantity one minus cotangent of x divided by quantity tangent of x minus one

OpenStudy (anonymous):

\[1 - \cot x \over \tan x -1\]

OpenStudy (anonymous):

we know: \[\cot x = \frac{ \cos x }{ \sin x }\] and \[\tan x = \frac{ \sin x }{ \cos x }\] next, change cot and tan with them. \[= \frac{ 1 - \frac{ \cos x }{ \sin x } }{ \frac{ \sin x }{ \cos x } -1 }\] \[ 1 - \frac{ \cos x }{ \sin x } = \frac{ \sin x - \cos x }{ \sin x } \] \[{\frac{ \sin x }{ \cos x } -1} = \frac{ \sin x - \cos x }{ \cos x } \] so, \[= \frac{ \frac{ \sin x - \cos x }{ \sin x } }{ \frac{ \sin x - \cos x }{ \cos x } }\] \[=\frac{ \frac{ 1 }{ \sin x } }{ \frac{ 1 }{ \cos x } }\] \[= \frac{ \cos x }{ \sin x }\] \[= ....\] hehe :)

OpenStudy (anonymous):

how did you get 1= cosx/sinx ?

OpenStudy (anonymous):

I meant 1 -*

OpenStudy (anonymous):

oh okay, you know \[1- \frac{ 2 }{ 5 }\] is same with \[\frac{ 5 }{ 5 } - \frac{ 2 }{ 5 } = \frac{ 3 }{ 5 }\]? is same with \[1-\frac{ \sin x }{ \cos x }\] so 1 is same with \[\frac{ \cos x }{ \cos x }\] so we can calculate to \[\frac{ \cos x }{ \cos x } - \frac{ \sin x }{ \cos x }\] it will be same with \[\frac{ \cos x - \sin x }{ \cos x }\] right? hehehe :)

OpenStudy (anonymous):

in my answers is \[1 - \frac{ \cos x }{ \sin x }\] it mean 1 is same with \[\frac{ \sin x }{ sin x }\] so, \[\frac{ \sin x }{ \sin x } - \frac{ \cos x }{ \sin x } = \frac{ \sin x - \cos x }{ \sin x }\]

OpenStudy (anonymous):

Wow thank you, lol. Can you help me with two more problems? \[\tan \theta \over \cot \theta \] I got to \[\frac{ \sin }{ \cos } \over \frac{ \cos }{ \sin }\] and then I am kind of stuck from here

OpenStudy (anonymous):

first, thanks for medal :) of course,hehe i think the answers is 1. we know cot is 1/tan x. so \[\tan x \frac{ 1 }{ \tan x } = 1\]

OpenStudy (anonymous):

Thanks and your welcome

OpenStudy (anonymous):

sorry i was make mistake with second question

OpenStudy (anonymous):

the answers is tan^2 x

OpenStudy (anonymous):

it must \[tanx \frac{ 1 }{ cotx }\] \[tanx \frac{ 1 }{ \frac{ 1 }{ tanx } }\] so, tanx go up, then multiply tanx with tanx = \[\tan^{2}x\]